The orbital speed of the Earth around the Sun is 30 km/s. In one year, how many seconds do the clocks on the Earth lose with respect to the clocks of an inertial reference frame at rest relative to the Sun


1
Expert's answer
2021-11-16T19:20:39-0500

If we are to consider an inertial frame at rest relative to the sun, it must be the one keeping the proper time interval ∆tp. Over the course of one year, we have

"∆t_p = 1 \; yr ≈ 3.156 \times 10^7 \;s"

The observers on earth are in motion relative to the sun, and therefore they measure a dilated time

interval "\u2206t^* =\u03b3\u2206t_p" . We are asked for the difference between the two clocks after one year as measured in the sun’s inertial frame, or

"difference = \u2206t^* \u2212 \u2206t_p = \u03b3\u2206t_p \u2212 \u2206t_p = (\u03b3 \u2212 1) \u2206t_p"

Since in this case the relative velocity of earth is small with respect to c, v = 30 km/s = 3 × 104 s so v/c= 10−4 , we can use the second approximation given.

"γ -1 = \frac{1}{\sqrt{1 -\frac{v^2}{c^2}}} -1 ≈ 1 + \frac{1}{2} \frac{v^2}{c^2} -1 = \frac{1}{2} \frac{v^2}{c^2 } \\

difference ≈ \frac{1}{2} \frac{v^2}{c^2} Δt_p \\

= \frac{1}{2} (10^{-4})^2 (3.156 \times 10^7 \;s) \\

= (5 \times 10^{-9}) (3.156 \times 10^7 \; s) \\

= 0.16 \;s"